OpenStudy (anonymous):

linear equations. gradient/slope is 1.8. 2 co-ordinates, (100, 212) and (0, 32). Where does the y incercept the line?

OpenStudy (anonymous):

y = mx + b. y = 180 + b 212 = 180 + b 212 - 180 = b 32 = b right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ahh, good, thankyou